A solution of carbon tetrachloride and acetic acid CH3COOH is boiling at 87°C. A sample of the vapor above the solution is cooled until it condenses. This condensed sample is analyzed, and turns out to be 90.% carbon tetrachloride by mass.

Calculate the percent by mass of carbon tetrachloride in the boiling solution. Here's some data you may need:

normal boiling point density vapor pressure at

87°C

carbon tetrachloride

77.°C

1.6gmL

1011.torr

acetic acid

118.°C

1.1gmL

406.torr

Be sure your answer has the correct number of significant digits.

Note for advanced students: you may assume the solution and vapor above it are ideal.
Answer #1

Given that In vapour phase Mass % of CCl4 = 90%

CCl4 = A and acetic acid = B
let the mass of CCl4 be mA and that of acetic acid be mB in vapour phase
Given: mA = 0.90 * (mA+mB)
mA = 9*mB
molar mass of CCl4 = 154 g/mol
molar mass of acetic acid = 60 g/mol
In vapour phase, as number of moles of CCl4, nA = mA/molar mass of CCl4 = mA/154
similarly, number of moles of acetic acid, nB = mB/molar mass of acetic acid= mB/60
or nB = mA/540
thus nA+nB = mA (1/154 + 1/540)
mole fraction of CCl4 in vapour phase, yA = nA/(nA+nB)
upon substitution we get: yA = (mA/154)/(mA/154 + mA/540)
yA = 0.78
Total pressure in vapour phase = pressure exerted by each fraction
PT = PA + PB
From Raoult's Law PA = xA. PA0 and PB = xB. PB0 where xA and xB are the mole fractions of the components in solution
where xA+xB = 1
upon substitution we get: PT = PB0 + (PA0-PB0)xA
using PA0 = 1011 torr and PB0 = 406 torr
PT = 406 + (406-1011)xA = 406 - 605 xA
and from Daltons's Law PA = yA.PT
thus xA. PA0 = yA.PT
xA . 1011 = 0.78 *(406 - 605 xA)
xA = 0.21 and xB = 1-xA = 0.79
using xA/xB = mAMB/mBMA
0.21/0.79 = 60*mA/154*mB
mB/mA = 1.466
add 1 both sides and then inverse
mA+mB/mA = 2.466
or mA/(mA+mB) *100 = 40.55%
thus in solution mass percent of CCl4 is 40.5 %